# Vision-Language Models can Identify Distracted Driver Behavior from Naturalistic Videos

### v1

@@ III-A Overview of the Proposed Frameworks
We fundamentally propose two methodologies: (i) frame-based CLIP models (Zero-shotCLIP, Single-frameCLIP, and Multi-frameCLIP) and (ii) a video-based CLIP model (VideoCLIP). An overview of these frameworks is provided in Fig. 1. CLIP [4] is a popular vision-language model. It is proficient in most visual understanding tasks, since it was trained on a large variety (around 400 million) of image-text pairs from the web.
@@
Figure 2: The proposed multimodal Single-frameCLIP framework that exploits the pretrained CLIP architecture to extract vision-text semantic information.
@@
The frame-based models operate by taking inputs similar to the original CLIP model and making predictions on distracted actions. As shown in Fig. 2, the model projects the input image and text pairs into a multi-modal latent space and minimizes a contrastive loss, which enables it to learn meaningful and useful visual and textual representations that can be leveraged across different tasks. Therefore, the base model can be fine-tuned for task-specific learning. Given its ability to be instructed by natural language for a broad range of downstream tasks, such as video question answering and understanding actions from video, we develop our distracted driving analysis framework based on this model. A great advantage of using a pre-trained network over other strategies is the ability to utilize the pre-existing visual representation of the language prompt, offering the zero-shot transfer. In this work, we initially leverage CLIP’s pre-trained visual embedding for zero-shot transfer on distracted driving datasets, categorizing the distracted action classes without any additional training. Following this, we fine-tune a linear classifier layer built upon the original pre-trained CLIP model. The yellow block shown in Fig. 2 represents the frozen layers. We add the green block positioned on top of the yellow layers, which represents the trainable layer. This task-based, fine-tuning green block involves training a straightforward classifier layer using the pre-existing embeddings from the frozen layers for a specific task. In our case, this task is to classify distracted driving actions from visual data. We evaluate the performance of the whole framework on a corresponding distracted driving dataset. As per the primary model [4], we implement five CLIP visual backbones- ViT-L/14, ViT-B/16, ViT-B/32, RN101 and assess each of their performance individually.
The Zero-shotCLIP, Single-frameCLIP and Multi-frameCLIP frameworks have input $I\in\mathbb{R}^{3\times H\times W}$ where $I$, $H$, and $W$ represent the input image, image height and image width, respectively. The Zero-shotCLIP makes predictions on raw frames without requiring any training. The key difference between the other two frameworks is that Single-frameCLIP makes predictions looking at a single frame and the Multi-frameCLIP makes predictions on a stack of frames conducting a majority vote at the classification layer. For both frame-based and video-based CLIP all the layers up to the classifier layer were kept frozen and we only finetuned the last trainable layer (shown in Fig. 2).
@@
Fig. 3 shows the overall sampling procedure that we follow for our experiments. The blue block on the left shows some sample frames from the original video, which has a frame rate of 30 frames per second (FPS). The yellow blocks show the sampled frames at corresponding sampling rates. Due to the CLIP network’s limitations in processing numerous frames simultaneously, we sample the frames from the video data at various sampling rates and created a 30-frame-long video piece to feed into the VideoCLIP network. From Fig. 3 , it could be observed that when the sampling rate is high, such as 20 FPS, the consecutive frames exhibit little semantic differences, and the video snippet (built from 30 frames) covers less time duration. In contrast, if the sampling rate is low, the video snippet could span longer time duration, with greater semantic differences between two consecutive frames. We perform a series of experiments to determine the best sampling rate for the video-based framework, discussed in Section IV, and the results of the experiments are explained in Section V.
Figure 3: The raw data is shown in the blue box, which contains 300 frames of an event lasting for 10 seconds. In other words, each second covers 30 frames. We sample frames with different sampling rates from the raw data, and the sampled data is shown in the yellow box. For example, when the sampling rate is 2 frames per second (FPS), we sample two images from 30 raw frames covered in each second. Since the raw data lasts for 10 seconds, we will have a total of 20 sampled images in the end.
@@ IV-A Dataset
We evaluate the proposed frameworks on two commonly used distracted driving datasets, namely SynDD1 [38] and StateFarm [39]. Both of these datasets serve as standard benchmarks for the analysis of distracted driving behavior, typically providing annotated images and videos that correspond to diverse distracted driving behaviors exhibited by participants. Each dataset comes with an individual driver ID for each participant. The selection of these datasets was motivated by the availability of unique driver IDs for each participant, enabling us to conduct ‘leave a set of drivers out’ experiments, ensuring the absolute separation of testing and training sets. For our experiments, we utilize the driver IDs to separate out a specific set of drivers only for testing. In addition, we assess the performance of our distracted driving models through a $k$-fold cross-validation approach, executed on each driver set. Then we calculate the average accuracy across all $k$ folds and report it as the final accuracy of the frameworks. Moreover, we analyze the proposed frameworks’ performances for each dataset separately and present the resulting analyses in detail in the subsequent sections.
@@ IV-B SynDD1 dataset
The SynDD1 dataset [38] is a recent addition to distracted driving datasets, offering insights into naturalistic distracted driving behaviors. The dataset includes videos of distracted driving events, recorded by cabin cameras inside the vehicle. A total of 18 participants performed various in-car driving activities, which were recorded (as video) from three distinct camera angles - the dashboard camera, the rear-view mirror camera, and the right-side window camera. The videos have two types of annotation files - gaze activity and distracted activity. The first row of Fig. 4 shows some sample frames of distracted driving actions from the dashboard camera view. There are 18 driving activities performed by the participants including ‘normal forward driving’, ‘drinking’, ‘phone call (right hand)’, ‘phone call (left hand)’, ‘eating’, ‘texting (right hand)’, ‘texting (left hand)’, ‘hair and makeup’, ‘reaching behind’, ‘adjusting control panel’, ‘pick up something from the floor (driver seat)’, ‘pick up something from the floor (passenger seat)’, ‘talking to the passenger at right’, ‘talking to the passenger at the backseat’, ‘yawning’, ‘hand on head’, ‘singing’, ‘shaking or dancing’. Each video file comes with annotations, including action types, the start and end time of an action, the appearance of the participant and so on. Therefore, we have the start and end time of a particular distracted action and their corresponding time duration. Each distracted driving video in the dataset has a frame rate of 30 FPS, with each frame possessing an RGB resolution of (1080,1920,3). For the purpose of this study, we consider the distracted behaviors of 14 drivers, thereby generating a more concise distracted driving dataset. For further generalization, we merge the similar action categories and consider eight distinct distracted actions shown in Table I. This modification allowed us to compare the results of our proposed frameworks for both datasets with an almost consistent number of classes. Fig. 5 shows the distribution of the distracted driving action duration (in seconds) across eight distinct distracted driving action classes derived from the SynDD1 dataset. The X-axis enumerates the eight specific distracted driving action categories, while the Y-axis represents the corresponding duration (sec) of these actions aggregated across all drivers. The average action duration across all categories is approximately 18.692 seconds. The median and spread of the action duration suggest an almost uniform distribution. Consequently, the probability of a specific distracted driving action being more prevalent than others within the examined range is nearly non-existent. The likelihood of occurrence for all durations of distracted driving actions is essentially equal. However, it is important to note that perfectly uniform distributions are rare to see in real-world scenarios, especially in studies related to human behavior.
@@
3 driver is picking something from floor while driving a car
@@ IV-C StateFarm Dataset
The StateFarm dataset [39] serves as another benchmark dataset for distracted driving behavior analysis, providing only image (RGB) data. The images were captured from an in-cabin camera positioned on the right-side window, offering a right-side perspective of the driver, as seen in the second row of Fig. 4. Data was collected from 26 participants, and the total number of image frames is 22,424. Similar to the SynDD1 dataset, the StateFarm dataset includes unique driver IDs for each participant, which enables us to perform driver-level k-fold cross-validation. However, it does not provide information on the action duration, start, or end times of actions as the labels pertain only to image frames. There are ten distracted driving activity categories in total- ‘normal driving’, ‘texting (right hand)’, ‘talking on the phone (right hand)’, ‘texting (left hand)’, ‘talking on the phone (left hand)’, ‘operating the radio’, ‘drinking’, ‘reaching behind’, ‘hair and makeup’, ‘talking to a passenger’. The images have a resolution of (640,480,3). We consider all the drivers and ten action categories for our experimental analysis.
IV-D Experiment Setup - ZeroshotCLIP and Single-frameCLIP
@@
Our overall approach is depicted in Fig. 2. First, we explore the zero-shot transfer capability of the CLIP model on distracted driving data in the Zero-shotCLIP framework. Then, we explore the task-based finetuning route to evaluate the “task learning capability” of our proposed Single-frameCLIP framework. For the Zero-shotCLIP experiment, there is no training associated, and we simply predict the class-level confidence score of the pre-trained CLIP model on raw image frames (224x224x3). Next, we assess the representation learning capability of the pre-trained CLIP model on the distracted driving dataset in the Single-frameCLIP framework. We add an extra linear classifier layer on top of the pre-trained CLIP model architecture and train the single layer on distracted driving datasets (SynDD1, StateFarm). For training the linear classifier (see CLIP [4]), we use image frames of input size (224x224x3) collected from the above two datasets. We utilize a batch size of 100 and a learning rate of 0.001. While training the Single-frameCLIP, we kept the pre-trained layers frozen and updated only the final layer based on the cross-entropy loss. Fig. 1 provides a visual representation of this approach, which generates predictions on a frame-by-frame basis. To illustrate this process, let’s consider a 10-second event sampled at a rate of 30 frames per second with a total of 300 frames (depicted in Fig. 1 by the yellow box). For each of these frames, the Single-frameCLIP model generates a prediction, beginning with the first frame (Image-1), and subsequently for each following frame (Image-2, Image-3, etc). In addition, we perform 7-fold and 8-fold cross-validation based on the driver ID for the SynDD1 and StateFarm datasets, and report the average accuracy in Section V. Moreover, we experiment with various CLIP backbones (see CLIP [4]) and other hyperparameters. Notably, the original CLIP model employs semantic information in text labels rather than the traditional one-hot labels, which require providing a textual description or prompt for each action category. Therefore, we utilized a full sentence as a prompt for describing the image frame of each action class instead of a single word. The prompt descriptions for the SynDD1 dataset are shown in Table I. As reported in [4], we consider manually fine-tuning and rephrasing the prompt sentences for our experiments, an area that can be further optimized in future work.
IV-E Experiment Setup - Multi-frameCLIP
@@
The Multi-frameCLIP framework isspecifically designed to analyze sequences of consecutive frames from distracted driving videos. This approach extends upon the Single-frameCLIP, concluding with an additional majority voting step applied across consecutive frames to generate the final event prediction. The majority voting method allows our framework to make a single prediction on a particular action event. During the first step of training, this framework is provided with the frames between the start and end times of a specific action within the video footage. Following this, the Single-frameCLIP model is applied to each frame falling within these timestamps, providing individual predictions for each frame. A majority voting method is then employed to produce a final prediction. If this prediction aligns with the ground truth action label for that segment of the video, it is considered a correct prediction. This entire process is visually represented in Fig. 1. From the raw video data, we sample frames at varying rates (as depicted in Fig. 3). Let’s consider a 10-second event, containing 300 frames captured at a rate of 30 FPS (shown by the yellow box). The Single-frameCLIP model is employed to make predictions for each of these frames, starting with the first frame (Image-1), and then sequentially progressing (Image-2, Image-3, and so on). Finally, a majority vote among all the individual frame predictions is taken to make a final prediction for the entire 10-second event in the video. This method allows Multi-frameCLIP to extend Single-frameCLIP’s capability by incorporating a temporal understanding of the events in a video. To ensure robust results, we conducted cross-validation experiments with sampling rates of 1, 2, 5, 20, and 0.5 frames per second.
IV-F Experiment Setup - VideoCLIP

### v3

@@ III-A Proposed Frameworks
We fundamentally propose two methodologies: (i) frame-based models (Zero-shotCLIP, Single-frameCLIP, and Multi-frameCLIP) and (ii) a video-based model (VideoCLIP). An overview of these frameworks is provided in Fig. 1. CLIP [10] is a popular vision-language model. It is proficient in most visual understanding tasks since it was trained on a large variety (around 400 million) of image-text pairs from the web.
@@
Figure 2: The proposed multimodal Single-frameCLIP framework exploits the pretrained CLIP architecture to extract vision-text semantic information.
@@
The frame-based models operate by taking inputs similar to the original CLIP model and making predictions on distracted actions. As shown in Fig. 2, the model projects the input image and text pairs into a multi-modal latent space and minimizes a contrastive loss, which enables it to learn meaningful and useful visual and textual representations. Therefore, the base model can be fine-tuned for task-specific learning. Given its ability to be instructed by natural language for a broad range of downstream tasks, such as video question answering and understanding actions from video, we develop our distracted driving analysis framework based on this model. A great advantage of using a pre-trained network over other strategies is the ability to utilize the pre-existing visual representation of the language prompt, offering the zero-shot transfer. In this work, we initially leverage CLIP’s pre-trained visual embedding for zero-shot transfer on distracted driving datasets and categorize the distracted action classes without any additional training. Following this, we fine-tune a linear classifier layer built upon the original pre-trained CLIP model. The yellow block shown in Fig. 2 represents the frozen layers. We add the green block positioned on top of the yellow layers, which represents the trainable layer. This task-based fine-tuning green block involves training a straightforward classifier layer using the pre-existing embeddings from the frozen layers for a specific task. In our case, this task is to classify distracted driving actions from visual data. We evaluate the performance of the whole framework on corresponding distracted driving datasets. As per the primary model [10], we implement six CLIP visual backbones- ViT-L/14, ViT-L/14@336px, ViT-B/16, ViT-B/32, RN101 and assess each of their performance individually.
The Zero-shotCLIP, Single-frameCLIP and Multi-frameCLIP frameworks have input $I\in\mathbb{R}^{3\times H\times W}$ where $I$, $H$, and $W$ represent the input image, image height and image width, respectively. The Zero-shotCLIP makes predictions on raw frames without requiring any training. The key difference between the other two frameworks is that Single-frameCLIP makes predictions looking at a single frame and the Multi-frameCLIP makes predictions on a stack of frames conducting a majority vote at the classification layer. For both frame-based and video-based CLIP, all the layers up to the classifier layer were kept frozen and we only finetuned the last trainable (green) layer (shown in Fig. 2).
@@
Fig. 3 shows the overall sampling procedure that we follow for our experiments. The blue block on the left shows some raw frames from the original video, which has a frame rate of 30 frames per second (FPS). The yellow blocks show the sampled frames at corresponding sampling rates. Due to the CLIP network’s limitations in processing numerous frames simultaneously, we sample the frames from the video data at various sampling rates and create a 30-frame-long video clip to feed into the VideoCLIP network. From Fig. 3, it could be observed that when the sampling rate is high, such as 20 FPS, the consecutive frames exhibit little semantic differences, and the video snippet (created from the 30 frames) covers less time duration. In contrast, if the sampling rate is low, the video snippet could span a longer time duration, with greater semantic differences between two consecutive frames. We perform a series of experiments to determine the best sampling rate for the video-based framework, discussed in Section IV, and the results of the experiments are explained in Section V.
Figure 3: The raw data is shown in the blue box, which contains 300 frames of an event lasting for 10 seconds. In other words, each second covers 30 frames. We sample frames with different sampling rates from the raw data, and the sampled data is shown in the yellow box. For example, when the sampling rate is two frames per second (FPS), we sample two images from 30 raw frames covered in each second. Since the raw data lasts for 10 seconds, we will have a total of 20 sampled images in the end.
@@ IV-A Dataset
The datasets used in our study are specifically chosen for their unique driver IDs or folders for each participant. This feature is crucial for conducting our ‘leave a set of driver out’ approach while testing. We evaluate our proposed frameworks on four distinct distracted driving datasets - DMD [35], SAM-DD [4], SynDD1 [65] and StateFarm [66]. These datasets serve as standard benchmarks for the analysis of distracted driving behavior, typically providing annotated images and videos that correspond to diverse distracted driving behaviors exhibited by participants. Each dataset comes with an individual driverID (SAM-DD has separate folders for each participant) for each participant.
For our experiments, we utilize the driverIDs to separate out a specific set of drivers only for testing purposes, ensuring no overlapping between training and testing splits. Therefore, our framework does not get the chance to memorize the driver’s appearance and only focuses on distracted actions. We believe this is crucial for properly validating our algorithm’s effectiveness. This approach addresses the inaccuracies encountered in previous studies where such separation was not maintained. Moreover, to ensure consistency and reliability in our experimental approach, we executed all our experiments for all the proposed models with the same experimental settings across all the datasets we used. In addition, we assess the performance of our distracted driving models through a $k$-fold cross-validation approach executed on each driver set. Then, we calculate the average accuracy across all the $k$ folds and report it as the final accuracy of the models. Moreover, we analyze the proposed frameworks’ performances for each dataset separately and report them in the subsequent sections.
@@ IV-B SynDD1 Dataset
The SynDD1 dataset [65] is a recent addition to distracted driving datasets, offering insights into naturalistic distracted driving behaviors. The dataset includes videos of distracted driving events recorded by cabin cameras inside the vehicle. A total of 18 participants performed various in-car driving activities, which were recorded (as video) from three distinct camera angles - the dashboard camera, the rear-view mirror camera, and the right-side window camera. The videos have two types of annotation files - gaze activity and distracted activity. The first row of Fig. 4 shows some sample frames of distracted driving actions from the dashboard camera view. There are 18 driving activities performed by the participants, including ‘normal forward driving’, ‘drinking’, ‘phone call (right hand)’, ‘phone call (left hand)’, ‘eating’, ‘texting (right hand)’, ‘texting (left hand)’, ‘hair and makeup’, ‘reaching behind’, ‘adjusting control panel’, ‘pick up something from the floor (driver seat)’, ‘pick up something from the floor (passenger seat)’, ‘talking to the passenger at right’, ‘talking to the passenger at the backseat’, ‘yawning’, ‘hand on the head’, ‘singing’, ‘shaking or dancing’. Each video file comes with annotations, including action types, the start and end time of an action, the appearance of the participant and so on. Therefore, we have the start and end time of a particular distracted action and their corresponding time duration. Each distracted driving video has a frame rate of 30 FPS. Each frame has an RGB resolution of 1920x1080 pixels. For the purpose of this study, we consider the distracted behaviors of 14 drivers, thereby generating a more concise distracted driving dataset. For further generalization, we merge the similar action categories and consider eight distinct distracted actions shown in Table I. Fig. 5 shows the distribution of the distracted driving action duration (in seconds) across eight classes. The X-axis enumerates the eight specific distracted driving action categories, while the Y-axis represents the corresponding duration (sec) of these actions aggregated across all drivers. The average action duration is approximately 18.692 seconds. The median and spread of the action duration suggest an almost uniform distribution. Consequently, the probability of a specific distracted driving action being more prevalent than others within the examined range is nearly zero. Therefore, the likelihood of occurrence for all durations of distracted driving actions is essentially equal. However, it is important to note that perfectly uniform distributions are rare to see in real-world scenarios, especially in studies related to human behavior.
@@
3 driver is picking something from the floor while driving a car
@@ IV-C StateFarm Dataset
The StateFarm dataset [66] serves as a common dataset for distracted driving behavior analysis, providing only image (RGB) data. The images were captured from an in-cabin camera positioned on the right-side window, offering a right-side perspective of the driver, as (d), (e) and (f) of Fig. 4. Data were collected from 26 participants, and the total number of image frames is 22,424. Similar to the SynDD1 dataset, the StateFarm dataset includes unique driver IDs for each participant, which enables us to perform the ‘leave a set of driver out’ step and driver-level k-fold cross-validation. However, it does not provide information on the action duration, start, or end times of actions as the labels pertain only to image frames. There are ten distracted driving activity categories in total- ‘normal driving’, ‘texting (right hand)’, ‘talking on the phone (right hand)’, ‘texting (left hand)’, ‘talking on the phone (left hand)’, ‘operating the radio’, ‘drinking’, ‘reaching behind’, ‘hair and makeup’, ‘talking to a passenger’. The images have a resolution of (640,480,3). We consider all the drivers and ten action categories for our experimental analysis.
IV-D DMD Dataset
The Driver Monitoring Dataset (DMD) is a specialized dataset  [35] designed to facilitate research in driver distraction and behavior recognition. This dataset captures various aspects of driver activity in 37 participants- focusing on distraction, drowsiness, hand movements, and gaze direction. This dataset offers RGB images (up to 1920x1080 pixels), depth data, and corresponding metadata with ten distraction categories. The data was recorded in two different environments - outdoors in a real car and indoors in a driving simulator. There are ten distracted driving actions, including ‘drinking water’, ‘adjusting hair and makeup’, ‘talking to the phone on right hand’, ‘adjusting radio’, ‘reaching to the backseat’, ‘reaching side’, ‘driving safely’, ‘talking to the passenger’, ‘texting on the phone with right hand’ and ‘yawning’. We implemented a 7-fold cross-validation on the ten distraction classes with 78917 samples for each fold and 552419 samples in total for all the folds. In addition, we performed the ‘leave a set of drivers out’ method using the provided driver IDs within the 7-fold cross-validation to evaluate our model performance.
IV-E SAM-DD Dataset
The Singapore AutoManNTU distracted driving (SAM-DD) dataset [4] is a comprehensive dataset for intelligent driving research. It includes image data from 42 participants (34 males, 8 females) of diverse ages and driving experiences. It features 51,175 samples across front and side camera views, encompassing both RGB and Depth image modes. The dataset includes ten distinct distraction classes with nine physically distracted behaviors - ‘drinking’, ‘talking’, ‘texting’, ‘touching hair’, ‘adjusting glasses’, ‘reaching behind’, and a fatigue-related behavior (head dropping). The high resolution (1200x900 pixels) RGB data makes it suitable for developing learning-based models in various driving environments and driver state-related applications. Both the DMD and SAM-DD are large datasets in terms of distracted class, number of samples per class and number of drivers. To ensure consistency and reliability in our experimental approach, we executed our proposed models with the same experimental settings for all the datasets.
IV-F Experiment Setup - ZeroshotCLIP and Single-frameCLIP
@@
Our overall approach is depicted in Fig. 2. First, we explore the zero-shot transfer capability of the CLIP model on distracted driving data in the Zero-shotCLIP framework. Then, we explore the task-based finetuning route to evaluate the “task learning capability” of our proposed Single-frameCLIP framework. For the Zero-shotCLIP experiment, there is no training associated, and we simply predict the class-level confidence score of the pre-trained CLIP model on raw image frames (224x224x3). Next, we assess the representation learning capability of the pre-trained CLIP model on the distracted driving datasets in the Single-frameCLIP framework. We add an extra linear classifier layer on top of the pre-trained CLIP model architecture and train the single layer on distracted driving datasets. For training the linear classifier (see CLIP [10]), we use image frames of input size (224x224x3) collected from the four datasets. We utilize a batch size of 100 and a learning rate of 0.001. While training the Single-frameCLIP, we kept the pre-trained layers frozen and updated only the final layer based on the cross-entropy loss. Fig. 1 provides a visual representation of this approach, which generates predictions on a frame-by-frame basis. To illustrate this process, let’s consider a 10-second event sampled at a rate of 30 frames per second with a total of 300 frames (depicted in Fig. 1 by the yellow box). For each of these frames, the Single-frameCLIP model generates a prediction, beginning with the first frame (Image-1) and subsequently for each following frame (Image-2, Image-3, etc). In addition, based on the driverID, we perform an 8-fold cross-validation for the StateFarm datasets and a 7-fold cross-validation for the other three datasets. We report the average accuracy in Section V. Moreover, we experiment with various CLIP backbones (see CLIP [10]) and other hyperparameters. Notably, the original CLIP model employs semantic information in text labels rather than the traditional one-hot labels, which require providing a textual description or prompt for each action category. Therefore, we provide a full sentence as a prompt for describing the image frame of each action class instead of a single word. The prompt descriptions for the SynDD1 dataset are shown in Table I. As reported in [10], we consider manually fine-tuning and rephrasing the prompt sentences for our experiments, an area that can be further optimized in future work.
IV-G Experiment Setup - Multi-frameCLIP
@@
The Multi-frameCLIP framework is specifically designed to analyze sequences of consecutive frames from distracted driving videos. This approach extends upon the Single-frameCLIP, concluding with an additional majority voting step applied across consecutive frames to generate the final event prediction. The majority voting method allows our framework to make a single prediction on a particular action event. During the first step of training, this framework is provided with the frames between the start and end times of a specific action within the video footage. Following this, the Single-frameCLIP model is applied to each frame falling within these timestamps, providing individual predictions for each frame. A majority voting method is then employed to produce a final prediction. If this prediction aligns with the ground truth action label for that segment of the video, it is considered a correct prediction. This entire process is visually represented in Fig. 1. From the raw video data, we sample frames at varying rates (as depicted in Fig. 3). Let’s consider a 10-second event containing 300 frames captured at a rate of 30 FPS (shown by the yellow box). The Single-frameCLIP model is employed to make predictions for each of these frames, starting with the first frame (Image-1) and then sequentially progressing (Image-2, Image-3, and so on). Finally, a majority vote among all the individual frame predictions is taken to make a final prediction for the entire 10-second event in the video. This method allows Multi-frameCLIP to extend Single-frameCLIP’s capability by incorporating a temporal understanding of the events in a video clip. To ensure robust results, we conducted cross-validation experiments with sampling rates of 1, 2, 5, 20, and 0.5 frames per second.
IV-H Experiment Setup - VideoCLIP